Fermi National Laboratory


Accelerator Update

Friday December 5
The day shift began with operations monitoring store 3083, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

At 9:36 AM, Operations received permission to send beam to Switchyard.

The Tevatron (TeV) suffered a sixteen-house quench at 10:36 AM. Just before the quench, Operations had pulled out the Collider Detector Facility (CDF) Japanese Pots. Unfortunately, one pot decided to run itself all the way in to the beam, which, at this point, seems to have caused the quench. The investigation continues.

Operators made the TeV, CDF, and D-zero (D0) ready for access at noon.

A Booster expert began aperture scans at 1:13 PM. This shut down beam to MiniBooNE for the hour that the scans lasted.

At 1:19 PM, RF technicians reported that a Main Injector (MI) RF station (MIRF13) had died. This repair will be made later.

Operations began sending shots of antiprotons from Pbar to the Recycler at 2:30 PM.

CDF experimenters came out of access at 3:46 PM.

At 5 PM, vacuum techs reported that the TeV vacuum at sector D4 was bad and that part of the cryo system would have to be warmed up to make repairs.

At 7 PM, Pbar experts went into a study period.

The Recycler successfully transferred antiprotons to MI at 9:26 PM.

Saturday December 6
The midnight shift began with TeV vacuum problems continuing, with Pbar conducting studies, and with MiniBooNE taking beam.

Pbar resumed stacking at 1 AM.

At 7:43 AM, operators made the Booster ready for access.

A Linac expert began a study period at 9:23 AM.

Earlier it was thought that there might be a problem with the TeV sector C1 cryo system. At 10:39 AM, operators accessed C1 and discovered a cryo leak on a conning tower. Both the Cryo Run Coordinator the vacuum techs believed the cryo system in that area would have to be warmed up to nitrogen temperature to make repairs. Here is what the operators saw:


You can see the tower down the tunnel.

Here's the Conning tower.

At 12:16 PM, the Linac expert completed his studies and then accessed the H- Source. He came out and Linac was ready to deliver beam by 1 PM.

A Booster expert began work on a kicker (B:MKS08) at 1:30 PM. He finished at 5:30 PM.

As of 2:30 PM, here was the TeV update:

  • The C19 spool piece (conning tower) required a warm up to resolder leads.
  • Techs began replacing the D49 collimator to fix the vacuum leak in this area.
At 4 PM Recycler reported that their vacuum leak investigations were going well and that they needed more time. They worked until 8 PM, but couldn't find a leak.

At 11:37 PM, the MiniBooNE interlocks were down due to a high airflow reading. An interlock tech was contacted.

Sunday December 7
The midnight shift began with TeV work continuing.

Pbar resumed stacking at 12:35 AM.

At 1:45 AM, the safety tech had MiniBooNE repaired and Operations began sending beam.

Cryo tech reported at 5:45 AM that the sector C1 cryo system Dewar emptied and a quadrupole magnet was beginning to warm up. Operators investigated and found a helium leak on the C19 conning tower. (This tower was repaired on Saturday, but obviously failed during cooldown.) They found a vapor cloud and liquid helium dripping from the spool piece. The Cryo Run Coordinator said that techs would have to replace this device.

At 10:30 AM, accelerator experts reported it would take a week to repair the TeV.

Pbar stopped stacking to begin an emittance growth study and to cool the stack.

At 3 PM, Operations began shooting antiprotons from Pbar, through MI, and into Recycler.

Pbar resumed stacking at 5:42 PM.

Switchyard studies began at 8:39 PM.

Monday December 8
The midnight shift began with the TeV sector C1 warm up continuing, with Pbar conducting studies, with Switchyard conducting studies, and with MiniBooNE taking beam.

Pbar resumed stacking at 1:11 AM.

Switchyard studies ended around 2 AM. They reported getting beam to F49.

Operations made the TeV, MI, CDF, and MiniBooNE ready for access at 6 AM.

Today's Plans
The TeV is off for a week due to cryo and vacuum problems. General machine maintenance will be allowed.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 12/8/2003   email Fermilab